# latexText page

### v1

@@ III On Adapting Nyström Preconditioners in Image Reconstruction
This section first describes IRM in Section III-A and then shows how RNP can be adapted to accelerate IRM for solving (3) efficiently. Because the weightings in IRM change at each iteration, computing RNP must be computationally inexpensive. We leverage the modern GPU computational platform and the PyTorch library to present an on-the-fly implementation for computing RNP at the end of this section. Additionally, we show how RNP can be incorporated with APG [59] to efficiently handle (4). Furthermore, we propose efficient approaches to solve the associated WPM to reduce the computational cost involved in applying RNP. We consider separately the $\ell_{p}-\ell_{q}$ and $\ell_{2}-\ell_{1,\phi}$ cases.
@@ III-A $\ell_{p}-\ell_{q}$ Image Reconstruction Acceleration
Following [29, Lemma 1], we represent the $p$th power of $r\in\mathbb{R}\setminus\{0\}$ with $20<p<$ 2 as
@@
where $a_{p}=\frac{p}{2-p}$ and $b_{p}=\frac{2^{\frac{2}{2-p}}}{(2-p)p^{\frac{2}{2-p}}}$ are two positive scalars. The minimizer of (12) is
@@
where $(\bm{\mathrm{A}}\bm{\mathrm{x}})_{m}$ and $y_{m}$ represent the $m$th element of $\bm{\mathrm{A}}\bm{\mathrm{x}}$, $\bm{\mathrm{y}}$, respectively. $(\bm{\mathrm{L}}\bm{\mathrm{x}})_{l}^{\zeta}$ denotes the $\zeta$th element in the $l$th group of $\bm{\mathrm{L}}\bm{\mathrm{x}}$. Here, we assume each group has the same number of elements and consider cases where each group represents a scalar or vector. Using (12), we rewrite (14) as
@@
where $v_{m}$ and $z_{l}^{\zeta}$ denote the $m$ and $(l-1)\pi+\zeta$th element of $\bm{\mathrm{v}}$ and $\bm{\mathrm{z}}$, respectively. By using an alternating minimization algorithm, IRM [16] solves (15) by successively updating $\bm{\mathrm{v}}$, $\bm{\mathrm{z}}$, and $\bm{\mathrm{x}}$. Algorithm 2 summarizes the IRM steps. Using (13), we derive closed-form solutions for $\bm{\mathrm{v}}^{k}$ and $\bm{\mathrm{z}}^{k}$ at steps 2 and 3 of Algorithm 2, i.e.,
$\begin{array}[]{rcl}\bm{\mathrm{v}}^{k}&=&\frac{p}{2}|(\bm{\mathrm{A}}\bm{\mathrm{x}}^{k})-\bm{\mathrm{y}}|^{p-2},\\[5.0pt] \bm{\mathrm{z}}^{k}&=&\frac{q}{2}|\bm{\mathrm{L}}\bm{\mathrm{x}}^{k}|^{q-2},\end{array}$ (16)
where the powers $(p-2)$ and $(q-2)$ are element-wise.
@@
Since the Hessian $\bm{\Phi}_{k}$ is a SPSD matrix, we use CG to solve (19). However, solving (19) is time-consuming when $\bm{\Phi}_{k}$ is ill-conditioned. we use (11) to construct a preconditioner for $\bm{\Phi}_{k}$ and apply the preconditioned CG (PCG) method to solve (19). Because $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{k}$ change at each iteration, computing $\bm{\mathrm{P}}$ in (11) must be computationally inexpensive. Section III-C discusses an on-the-fly implementation of RNP to address this computational issue. If $p,q=2$, then (3) is a quadratic minimization problem, allowing us to directly apply PCG with RNP. When $p=2,q\neq 1$, IRM is still applicable, but APG is more efficient as the following subsection discusses.
@@
2:     $\bm{\mathrm{v}}^{k}=\arg\min_{\bm{\mathrm{v}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{v}})$
3:     $\bm{\mathrm{z}}^{k}=\arg\min_{\bm{\mathrm{z}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{z}})$
@@ III-B $\ell_{2}-\ell_{1,\phi}$ Image Reconstruction Acceleration
APG, which only uses first-order information and achieves the optimal worst-case convergence rate of $\mathcal{O}(1/k^{2})$ [60], is a very appealing algorithm for addressing (4). To further accelerate the convergence of APG, we adapt RNP to APG, resulting in the weighted accelerated proximal gradient (WAPG) algorithm that is summarized in Algorithm 3. For any $\bm{\mathrm{x}}_{1},\bm{\mathrm{x}}_{2}\in\mathbb{R}^{N}$, the Lipschitz constant $L_{f}^{\bm{\mathrm{P}}}$ of $f$ with respect to the $\bm{\mathrm{P}}$-norm satisfies the following inequality
@@
If $f,g$ are convex, the worst-case convergence rate of the cost function sequence generated by Algorithm 3 is $\frac{2L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}}{(k+1)^{2}}$ [61, 10.7.5] where $\bm{\mathrm{x}}^{1}$ and $\bm{\mathrm{x}}^{*}$ are the initial and global values, respectively. Clearly, for $\bm{\mathrm{P}}=\bm{\mathrm{I}}$, WAPG simplifies to APG. By choosing a suitable $\bm{\mathrm{P}}$, one can expect a lower bound than $\bm{\mathrm{P}}=\bm{\mathrm{I}}$ yielding faster convergence. Indeed, our experimental results in CT reconstruction illustrate the faster convergence of WAPG with RNP.
@@
where $\bm{\mathrm{s}}^{k}=\bm{\mathrm{u}}^{k-1}-\alpha\bm{\mathrm{P}}^{-1}\nabla f(\bm{\mathrm{u}}^{k-1})$, $\bar{\lambda}=\alpha\lambda$, and $\alpha$ represents the stepsize. By using the fact that
@@
where $B_{\infty,\psi}$ represents the $\ell_{\infty,q}$ mixed unit-norm ball, i.e.,
$\begin{array}[]{rl}B_{\infty,\psi}=&\Big{\{}\mathcal{Q}=\{\bm{\mathrm{q}}_{1},\bm{\mathrm{q}}_{2},\ldots,\bm{\mathrm{q}}_{G}\}:\\ &\quad\|\bm{\mathrm{q}}_{l}\|_{\psi}\leq 1,~{}\forall l=1,\ldots,G\Big{\}},\end{array}$ (23)
and $\psi\geq 1$ satisfies $\frac{1}{\phi}+\frac{1}{\psi}=1$. Note that the inner product in (22) is applied to each component of $\mathcal{Q}$ and $\mathcal{V}$, where $\mathcal{V}=\bm{\mathrm{L}}\bm{\mathrm{x}}$. With these, we rewrite (21) as
@@
where $\bm{\mathrm{L}}^{\mathcal{T}}$ denotes the adjoint of $\bm{\mathrm{L}}$. After some rewriting, (24) becomes
@@
$-2\bar{\lambda}\bm{\mathrm{L}}\,\mathrm{prox}_{\delta_{\mathcal{C}}}^{\bm{\mathrm{P}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}),$ (28)
@@
Both (28) and (29) require computing $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ and $\bm{\mathrm{P}}^{-1}$, which can be computationally expensive for a general $\bm{\mathrm{P}}$. A natural way to choose $\bm{\mathrm{P}}$ is to approximate the Hessian $\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}$. By using RNP, $\bm{\mathrm{P}}$ (see (11)) can be rewritten as $\bm{\mathrm{I}}+\bar{\bm{\mathrm{U}}}\bar{\bm{\mathrm{U}}}^{\mathrm{T}}$ with $\bar{\bm{\mathrm{U}}}=\bm{\mathrm{U}}\sqrt{\frac{1}{\hat{s}_{K}+\mu}(\hat{\bm{\mathrm{S}}}+\mu\bm{\mathrm{I}})-\bm{\mathrm{I}}}$, where the square root here is applied to the diagonal elements of the inner item. Clearly, $\bm{\mathrm{P}}$ has the same structure as $\bm{\mathrm{W}}$ proposed in Theorem 1, so getting $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ reduces to addressing a nonlinear equation (8). Thus, the total computation for $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ can be dramatically reduced, as demonstrated by our numerical experiments in Section IV-B. Note that ${\hat{s}}_{K}$ is the $K$th eigenvalue of $\bm{\Phi}\left<\bm{\Omega}\right>$, so the last column in $\bar{\bm{\mathrm{U}}}$ will be a zero vector. In practice, we found that using $\sqrt{{\hat{s}}_{K}}$ instead of ${\hat{s}}_{K}$ resulted in faster convergence. So, in this paper, we use ${\hat{s}}_{K}\leftarrow\sqrt{{\hat{s}}_{K}}$ in (11). Since $\bm{\mathrm{U}}$ is an orthogonal matrix, we have $\sigma^{\bm{\mathrm{P}}}_{\mathrm{min}}=1$, so the Lipschitz constant of the cost function in (27) becomes $2\bar{\lambda}^{2}\|\bm{\mathrm{L}}\|^{2}$.
@@
III-C On-the-Fly Nyström Preconditioner Implementation
@@
On classical CPU computational platforms, a parallel scheme may be used to evaluate $\bm{\Phi}\bm{\Omega}$. However, a central server still must collect the results to formulate $\bm{\mathrm{Y}}$, making communication time a potential bottleneck. Modern machine learning libraries (e.g., PyTorch [62] and TensorFlow [63]) support a batch mode, allowing efficient parallel computation of $\bm{\Phi}\bm{\Omega}$. Additionally, these libraries are highly optimized for modern GPU computational platforms, and the batch mode is easy to implement with basic Python knowledge. These features enable us to compute RNP on-the-fly. Table IV presents the wall time of computing RNP for CT reconstruction with different acquisitions on CPU and GPU computational platforms, clearly showing the advantage of using GPU.
@@
Debluring/Super-Resolution (3) with $p\in(0,1]$ $q=1$nabla, TV
@@
IV Numerical Experiments
This section studies the performance of using RNP on three different image reconstruction tasks: image deblurring and super-resolution with impulsive noise, and CT. Following [16], we use a TV regularizer for image deblurring and super-resolution tasks. For CT reconstruction, we consider wavelet, TV, and HS norm regularizers. Table I presents the associated minimization problems and regularizers for different image tasks in this part. All experiments were implemented with PyTorch and run on an NVIDIA GeForce RTX 3090.

### v2

@@ III On Adapting Nyström Preconditioners in Image Reconstruction
For impulsive noise, a natural data-fidelity term is the $\ell_{p}$ quasi-norm with $p\leq 1$, allowing image reconstruction by solving the $\ell_{p}-\ell_{q}$ problem. When the noise is assumed to be i.i.d. Gaussian, we use the $\ell_{2}$ norm as the metric in the data-fidelity term, resulting in the $\ell_{2}-\ell_{1,\phi}$ problem. This section presents two baseline algorithms for solving $\ell_{p}-\ell_{q}$ and $\ell_{2}-\ell_{1,\phi}$, and then study how they can be accelerated using RNP.
Following we first describe IRM in Section III-A and then show how RNP can be adapted to accelerate IRM for solving (3) efficiently. Because the weightings in IRM change at each iteration, computing RNP must be computationally inexpensive. We leverage the modern GPU computational platform and the PyTorch library to present an on-the-fly implementation for computing RNP at the end of this section. Additionally, we show how RNP can be incorporated with APG [60] to efficiently handle (4). Furthermore, we propose efficient approaches to solve the associated WPM to reduce the computational cost involved in applying RNP. We consider separately the $\ell_{p}-\ell_{q}$ and $\ell_{2}-\ell_{1,\phi}$ cases.
@@ III-A $\ell_{p}-\ell_{q}$ Image Reconstruction Acceleration
Following [29, Lemma 1], we represent the $p$th power of $r\in\mathbb{R}\setminus\{0\}$ with $0<p<2$ as
@@
where $a_{p}=\frac{p}{2-p}$ and $b_{p}=\frac{2^{\frac{2}{2-p}}}{(2-p)\cdot p^{\frac{p}{2-p}}}$ are two positive scalars. The minimizer of (12) is
@@
[…]
@@
[…]
$\bm{\mathrm{v}}^{k}=\frac{p}{2}|(\bm{\mathrm{A}}\bm{\mathrm{x}}^{k})-\bm{\mathrm{y}}|^{p-2}~~\text{and}~~\bm{\mathrm{z}}^{k}=\frac{q}{2}|\bm{\mathrm{L}}\bm{\mathrm{x}}^{k}|^{q-2},$ (16)
where the powers (p2)(p-2) and (q2)(q-2) are element-wise.
@@
Since the Hessian $\bm{\Phi}_{k}$ is a SPSD matrix, we use CG to solve (19). However, solving (19) is time-consuming when $\bm{\Phi}_{k}$ is ill-conditioned. We use (11) to construct a preconditioner for $\bm{\Phi}_{k}$ and apply the preconditioned CG (PCG) method to solve (19). Because $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{k}$ change at each iteration [^1] Note that $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{k}$ are changed at each outer iteration but remain unchanged within CG iterations when solving (19)., computing $\bm{\mathrm{P}}$ in (11) must be computationally inexpensive. Section III-C discusses an on-the-fly implementation of RNP to address this computational issue. If $p,q=2$, then (3) is a quadratic minimization problem, allowing us to directly apply PCG with RNP. When $p=2$ and $q\neq 1$, IRM is still applicable, but APG is more efficient, as discussed in the following subsection.
@@
2:  $\bm{\mathrm{v}}^{k}=\arg\min_{\bm{\mathrm{v}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{v}},\bm{\mathrm{z}}^{k-1})$
3:  $\bm{\mathrm{z}}^{k}=\arg\min_{\bm{\mathrm{z}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{v}}^{k},\bm{\mathrm{z}})$
@@ III-B $\ell_{2}-\ell_{1,\phi}$ Image Reconstruction Acceleration
APG, which only uses first-order information and achieves the optimal worst-case convergence rate of $\mathcal{O}(1/k^{2})$ [61], is an appealing algorithm for addressing (4). To further accelerate the convergence of APG, we adapt RNP to APG, resulting in the weighted accelerated proximal gradient (WAPG) algorithm that is summarized in Algorithm 3. For any $\bm{\mathrm{x}}_{1},\bm{\mathrm{x}}_{2}\in\mathbb{R}^{N}$, the Lipschitz constant $L_{f}^{\bm{\mathrm{P}}}$ of $f$ with respect to the $\bm{\mathrm{P}}$-norm satisfies the following inequality
@@
If $f,g$ are convex, the worst-case convergence rate of the cost function sequence generated by Algorithm 3 is $\frac{2L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}}{(k+1)^{2}}$ [62, Chapter 10.7.5] where $\bm{\mathrm{x}}^{1}$ and $\bm{\mathrm{x}}^{*}$ are the initial and optimal values, respectively. Clearly, for $\bm{\mathrm{P}}=\bm{\mathrm{I}}$, WAPG simplifies to APG. By choosing a suitable $\bm{\mathrm{P}}$, one can expect a lower bound than $\bm{\mathrm{P}}=\bm{\mathrm{I}}$ yielding faster convergence. Indeed, our experimental results in CT reconstruction illustrate the faster convergence of WAPG with RNP, i.e., experiments in Sections IV-B and IV-C.
@@
where $\bm{\mathrm{s}}^{k}=\bm{\mathrm{u}}^{k-1}-\alpha\bm{\mathrm{P}}^{-1}\nabla f(\bm{\mathrm{u}}^{k-1})$ and $\bar{\lambda}=\alpha\lambda$. By using the dual representation
@@
where $B_{\infty,\psi}$ denotes the $\ell_{\infty,\psi}$ mixed unit-norm ball, i.e.,
$B_{\infty,\psi}=\Big\{\mathcal{Q}=\{\bm{\mathrm{q}}_{1},\bm{\mathrm{q}}_{2},\ldots,\bm{\mathrm{q}}_{G}\}:\|\bm{\mathrm{q}}_{l}\|_{\psi}\leq 1,~\forall l\Big\},$ (23)
and $\psi\geq 1$ satisfies $\tfrac{1}{\phi}+\tfrac{1}{\psi}=1$, we rewrite (21) as
@@
where $\bm{\mathrm{L}}^{\mathcal{T}}$ denotes the adjoint of $\bm{\mathrm{L}}$. Denote by $\mathcal{V}=\bm{\mathrm{L}}\bm{\mathrm{x}}$ with $\mathcal{V}=\{\bm{\mathrm{v}}_{1},\bm{\mathrm{v}}_{2},\ldots,\bm{\mathrm{v}}_{G}\}$. The inner product in (22) is defined as $\sum_{l=1}^{G}\langle\bm{\mathrm{q}}_{l},\bm{\mathrm{v}}_{l}\rangle$. Reorganizing (24), we get
@@
$2\bar{\lambda}\bm{\mathrm{L}}\,\mathrm{prox}_{\delta_{\mathcal{C}}}^{\bm{\mathrm{P}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}),$ (28)
@@
Both (28) and (29) require computing $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ and $\bm{\mathrm{P}}^{-1}$, which can be computationally expensive for a general $\bm{\mathrm{P}}$. A natural way to choose $\bm{\mathrm{P}}$ is to approximate the Hessian $\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}$. By using RNP, $\bm{\mathrm{P}}$ (see (11)) can be rewritten as $\bm{\mathrm{I}}+\bar{\bm{\mathrm{U}}}\bar{\bm{\mathrm{U}}}^{\mathrm{T}}$ with $\bar{\bm{\mathrm{U}}}=\bm{\mathrm{U}}\sqrt{\frac{1}{\hat{s}_{K}+\mu}(\hat{\bm{\mathrm{S}}}+\mu\bm{\mathrm{I}})-\bm{\mathrm{I}}}$, where the square root here is applied elementwise. Clearly, $\bm{\mathrm{P}}$ has the same structure as $\bm{\mathrm{W}}$ proposed in Theorem 1, so getting $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ reduces to addressing a nonlinear equation (8). Thus, the total computation for $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ can be dramatically reduced, as demonstrated by our numerical experiments in Section IV-B. Note that ${\hat{s}}_{K}$ is the $K$th eigenvalue of $\bm{\Phi}\left<\bm{\Omega}\right$>, so the last column in $\bar{\bm{\mathrm{U}}}$ will be a zero vector. In practice, we found that using $\sqrt{{\hat{s}}_{K}}$ instead of ${\hat{s}}_{K}$ resulted in faster convergence. So, in this paper, we use ${\hat{s}}_{K}\leftarrow\sqrt{{\hat{s}}_{K}}$ in (11). Since $\bm{\mathrm{U}}$ is an orthogonal matrix, we have $\sigma^{\bm{\mathrm{P}}}_{\mathrm{min}}=1$, so the Lipschitz constant of the cost function in (27) becomes $2\bar{\lambda}^{2}\|\bm{\mathrm{L}}\|^{2}$.
@@
III-C Implementation Details of On-the-Fly RNP Method
@@
On classical CPU computational platforms, a parallel scheme may be used to evaluate $\bm{\Phi}\bm{\Omega}$. However, a central server still must collect the results to formulate $\bm{\mathrm{Y}}$, making communication time a potential bottleneck. Modern machine learning libraries (e.g., PyTorch [63] and TensorFlow [64]) support a batch mode, allowing efficient parallel computation of $\bm{\Phi}\bm{\Omega}$. Additionally, these libraries are highly optimized for modern GPU computational platforms, and the batch mode is easy to implement with basic Python knowledge. These features enable us to compute RNP on-the-fly. Table II presents the wall time of computing RNP for CT reconstruction with different acquisitions on CPU and GPU computational platforms, clearly showing the advantage of using GPU. However, RNP requires storing an additional $K$ images. As a result, it may encounter memory limitations when handling high-dimensional images on a GPU platform. Therefore, developing a memory-efficient approach to apply RNP to high-dimensional images would be an interesting future direction.
Remark 1.
In this section, we introduce IRM (respectively, WAPG) for solving (3) (respectively, (4)) and discuss how RNP can be incorporated for acceleration. However, in some special cases, e.g., $p=2$ and $\mathcal{C}=\mathbb{R}^{N}$, (3) is identical to (4). This naturally raises the question of how to choose between IRM and WAPG. From our experiments, when both IRM and WAPG are applicable, we find that IRM converges faster than WAPG in terms of iterations. However, since IRM requires more computation per iteration than WAPG, we consistently observe that WAPG converges faster than IRM in terms of wall time. We validate this observation in Section IV-A4 and Section IV-B1. In this paper, we only consider linear inverse problems. However, we believe that RNP also has the potential to accelerate numerical solvers for nonlinear inverse problems. One example is the full-waveform inversion [65] in geophysics, where truncated Newton methods are among the appealing algorithms. However, one needs to solve a linear equation at each iteration, and the related coefficient matrix (i.e., the Hessian matrix) changes at every iteration. So using effective on-the-fly preconditioners could significantly accelerate the numerical efficiency. We leave the investigation of this direction for future work, as it is beyond the scope of this paper.
IV Numerical Experiments
@@
Debluring/Super-Resolution (3) with $p\in(0,1]$ $q=1$ $\nabla$, TV
@@
TABLE II: Comparison of wall time for obtaining RNP in computed tomography reconstruction on CPU and GPU.
Parallel-beam Fan-beam
Wavelet TV HS Wavelet TV HS
$K$ 2020 2020 $100$ 2020 2020 $100$
GPU $0.11$s $0.10$s $0.40$s $0.11$s $0.09$s $0.46$s
CPU $0.48$s $0.40$s $4.20$s $0.48$s $0.52$s $4.25$s
This section studies the performance of using RNP on three different image reconstruction tasks: image deblurring and super-resolution with impulsive noise, and CT. Following [16], we use a TV regularizer for image deblurring and super-resolution tasks. For CT reconstruction, we consider wavelet, TV, and HS norm regularizers. Table I presents the associated minimization problems and regularizers for different image tasks in this part. The trade-off parameter $\lambda$ is selected through a grid search to maximize the PSNR. All experiments were implemented with PyTorch and run on an NVIDIA GeForce RTX 3090.